a car starts at rest covering a distance of 200m. It continues moving with a constant velocity for another 300m before coming to rest in the last 100m . Determine the velocity of the car when it was moving with a constant velocity


1
Expert's answer
2021-02-22T10:25:09-0500

Average speed:

"v=\frac{s}{t}=\frac{200+300+100}{t}=\frac{600}{t}"